### v2

# Uniqueness of invariant measures as a structural property of Markov kernels

Abstract
We identify indecomposability as a key measure–theoretic mechanism underlying uniqueness of invariant probability measures for discrete–time Markov kernels on general state spaces. The argument relies on the mutual singularity of distinct invariant ergodic measures and on the observation that uniqueness follows whenever all invariant probability measures are forced to charge a common reference measure.
Once existence of invariant probability measures is established, indecomposability alone is sufficient to rule out multiplicity. On standard Borel spaces, this viewpoint is consistent with the classical theory: irreducibility appears as a convenient sufficient condition ensuring indecomposability, rather than as a structural requirement for uniqueness.
The resulting proofs are purely measure–theoretic and do not rely on recurrence, regeneration, return–time estimates, or regularity assumptions on the transition kernel.
1 Introduction
The existence and uniqueness of invariant probability measures are central questions in the study of Markov chains and stochastic dynamical systems. Existence is commonly obtained through compactness or tightness arguments, often supported by Lyapunov–type drift conditions. Uniqueness, by contrast, is most often derived from stronger dynamical assumptions, such as Harris recurrence, regeneration techniques, or explicit control of return times to petite or small sets. These ideas form the backbone of the general theory developed by Meyn and Tweedie [9].
While extremely powerful, recurrence–based approaches intertwine existence, uniqueness and ergodic convergence within a single framework. In particular, uniqueness of the invariant probability measure is typically obtained as a consequence of positive Harris recurrence together with irreducibility or minorization conditions. As a result, uniqueness is often perceived as a dynamical property, intrinsically linked to long–term return behaviour. This paradigm is well illustrated both in the classical literature and in more recent developments, where Harris–type arguments are refined or revisited in various directions (see, for instance, Hairer and Mattingly [8], Douc, Fort and Guillin [5]).
However, in several important classes of models—such as non–Feller Markov chains, discontinuous random dynamical systems, or nonlinear time series—verifying Harris recurrence may be technically delicate or may require strong topological assumptions. This difficulty is well documented in the literature on random iterative models and stochastic difference equations, where the transition kernel may fail to be continuous and discontinuities may occur on sets of positive Lebesgue measure (see, for example, Attali [2], Duflo [6], and references therein). In many such situations, existence of an invariant probability measure can nevertheless be established by stability or Lyapunov–type arguments, independently of any recurrence property.
This observation naturally raises the following question: does uniqueness of the invariant probability measure genuinely rely on recurrence, or is it instead governed by more elementary structural properties of the Markov kernel? From a measure–theoretic viewpoint, the fundamental obstruction to uniqueness is the presence of nontrivial invariant measurable sets. This idea already appears, at least implicitly, in early work of Breiman [3, 4], in the context of ergodic theorems and laws of large numbers for Markov chains. In the classical theory, however, this structural viewpoint remains tightly interwoven with dynamical assumptions ensuring ergodic convergence.
The purpose of the present work is to isolate uniqueness as a purely structural problem. Our main contribution is to identify indecomposability as a purely structural condition that prevents nontrivial invariant measurable decompositions and ensures uniqueness of invariant probability measures once existence is known. Indecomposability is defined solely in terms of absorbing measurable sets and does not involve any recurrence, accessibility, or return–time assumption.
On standard Borel spaces, indecomposability is strictly weaker than classical irreducibility in general. The fact that the two notions coincide a posteriori once existence of an invariant probability measure is guaranteed should therefore be read as a clarification of the logical structure: irreducibility is not an additional assumption required for uniqueness, but a convenient and well–understood sufficient condition ensuring indecomposability in classical settings.
Once existence of invariant probability measures is known, indecomposability alone is sufficient to rule out the coexistence of several invariant probability measures. More precisely, if uniqueness fails, then the state space admits a nontrivial absorbing measurable subset. The proof relies on the mutual singularity of distinct invariant ergodic measures and on the observation that uniqueness follows whenever all invariant probability measures are forced to charge a common reference measure. This mechanism is entirely measure–theoretic and does not involve return–time estimates, regeneration schemes, or regularity assumptions on the transition kernel.
From this perspective, classical irreducibility assumptions do not constitute a dynamical requirement for uniqueness, but rather provide a transparent way of making the underlying structural mechanism explicit. Under irreducibility, this mechanism can be exhibited by introducing a resolvent–type kernel, defined as a convex combination of the iterates of the original transition kernel. The resulting one–step positivity property with respect to a $\sigma$–finite reference measure forces all invariant probability measures to charge the same measure and therefore enforces uniqueness.
Since every invariant probability measure of the original kernel is also invariant for the resolvent kernel, uniqueness immediately transfers whenever existence has been established by independent means. In particular, in the classes of models considered in [2], where existence follows from stability arguments rather than recurrence, uniqueness becomes a direct consequence of the present results.
The aim of this work is therefore not to strengthen classical ergodic theorems, but to clarify the logical structure underlying uniqueness. From this perspective, recurrence and regeneration emerge as genuinely dynamical properties governing convergence and long–term behaviour, while uniqueness appears as a structural consequence of indecomposability, revealed through a one–step positivity mechanism acting at the level of invariant measures.
2 Main result
Let $(E,\mathcal{B})$ be a standard Borel space and let $P$ be a discrete-time Markov kernel on $(E,\mathcal{B})$.
We introduce the structural notion that lies at the core of the uniqueness mechanism.
Definition 1 (Indecomposability).
Let $(E,\mathcal{B})$ be a measurable space and let $P$ be a Markov kernel on $(E,\mathcal{B})$. The kernel $P$ is said to be indecomposable if there exist no two disjoint nonempty measurable sets $A,B\in\mathcal{B}$ such that
[…]
Indecomposability isolates a purely structural obstruction to the uniqueness of invariant probability measures. Classical irreducibility assumptions are not required in what follows; they will only be invoked later as convenient sufficient conditions ensuring indecomposability in standard settings.
Remark 2.
Throughout the paper, the term “absorbing” refers to invariance in the forward sense only, i.e. $x\in A\Rightarrow P(x,A)=1$. No backward or symmetric invariance is assumed.
We first recall a fundamental structural property of invariant ergodic measures. Although the following result can be derived from the ergodic decomposition theorem under standard assumptions, we include a direct proof since our goal is to isolate the measure–theoretic core of the uniqueness mechanism, without relying on convex–analytic or topological arguments on the space of invariant measures. In particular, an observation underlying all subsequent arguments is that distinct invariant ergodic probability measures must be mutually singular.
Lemma 3.
Let $P$ be a Markov kernel on a measurable space $(E,\mathcal{E})$. Any two distinct PP–invariant ergodic probability measures are mutually singular.
Proof.
Let $\mu$ and $\nu$ be two distinct PP–invariant ergodic probability measures and set
$\eta:=\tfrac{1}{2}(\mu+\nu).$
Then $\eta$ is PP–invariant and $\mu\ll\eta$. Let $f:=d\mu/d\eta$. Since $\mu+\nu=2\eta$, we also have $d\nu/d\eta=2-f$, hence $0\leq f\leq 2$ $\eta$–a.s.
Step 1: $Pf=f$ $\eta$–a.s. Fix a bounded measurable function $\varphi\in L^{\infty}(E)$. Using $\mu=f\,\eta$ and the invariance of $\mu$, we get
$\int_{E}\varphi(x)\,f(x)\,\eta(dx)=\int_{E}\varphi(x)\,\mu(dx)=\int_{E}(P\varphi)(x)\,\mu(dx)=\int_{E}(P\varphi)(x)\,f(x)\,\eta(dx).$ (1)
We now rewrite the last term. Define the finite measure $\Pi$ on $(E\times E,\mathcal{E}\otimes\mathcal{E})$ by
$\Pi(dx,dy):=\eta(dx)\,P(x,dy).$
Since $\varphi$ is bounded and $f\in L^{1}(\eta)$ with $f\geq 0$, the function
$(x,y)\longmapsto|\varphi(y)|\,f(x)$
is $\Pi$–integrable:
$\int_{E\times E}|\varphi(y)|\,f(x)\,\Pi(dx,dy)=\int_{E}f(x)\left(\int_{E}|\varphi(y)|\,P(x,dy)\right)\eta(dx)\leq\|\varphi\|_{\infty}\int_{E}f\,d\eta<\infty.$
Hence Tonelli’s theorem applies and yields
$\int_{E}(P\varphi)(x)\,f(x)\,\eta(dx)$ $=\int_{E}\left(\int_{E}\varphi(y)\,P(x,dy)\right)f(x)\,\eta(dx)$
$=\int_{E\times E}\varphi(y)\,f(x)\,\Pi(dx,dy)$
$=\int_{E}\varphi(y)\left(\int_{E}f(x)\,P(x,dy)\,\eta(dx)\right).$
By definition of the action of $P$ on $L^{1}(\eta)$ (equivalently, by the duality $\int(P\varphi)\,f\,d\eta=\int\varphi\,Pf\,d\eta$), the last expression equals $\int_{E}\varphi(y)\,(Pf)(y)\,\eta(dy)$. Plugging this into (1) gives
$\int_{E}\varphi\,f\,d\eta=\int_{E}\varphi\,Pf\,d\eta\qquad\forall\varphi\in L^{\infty}(E),$
hence $Pf=f$ $\eta$–a.s.
Step 2: $f$ is constant $\mu$–a.s. and $\nu$–a.s. Since $Pf=f$ $\mu$–a.s. and $0\leq f\leq 2$, ergodicity of $\mu$ implies that $f$ is $\mu$–a.s. constant: $f=c$ $\mu$–a.s. for some $c\in[0,2]$. Similarly, $f=c^{\prime}$ $\nu$–a.s. for some $c^{\prime}\in[0,2]$.
If $c=c^{\prime}$, then $f$ is $\eta$–a.s. constant, hence $\mu=f\eta$ and $\nu=(2-f)\eta$ are both proportional to $\eta$; since they are probability measures, this forces $\mu=\nu$, a contradiction. Therefore $c\neq c^{\prime}$.
Step 3: the only possible constants are $c=2$ and $c^{\prime}=0$. Since $\mu$ is supported on $\{f=c\}$, we have
$1=\mu(E)=\int_{\{f=c\}}f\,d\eta=c\,\eta(f=c),\qquad\text{so}\qquad\eta(f=c)=\frac{1}{c}.$
Likewise, $\nu$ is supported on $\{f=c^{\prime}\}$ and $d\nu=(2-f)\,d\eta$, hence
$1=\nu(E)=\int_{\{f=c^{\prime}\}}(2-f)\,d\eta=(2-c^{\prime})\,\eta(f=c^{\prime}),\qquad\text{so}\qquad\eta(f=c^{\prime})=\frac{1}{2-c^{\prime}}.$
The sets $\{f=c\}$ and $\{f=c^{\prime}\}$ are disjoint (since $c\neq c^{\prime}$), thus
$1=\eta(E)\geq\eta(f=c)+\eta(f=c^{\prime})=\frac{1}{c}+\frac{1}{2-c^{\prime}}.$
If $c<2$ then $1/c>1/2$, and if $c^{\prime}>0$ then $1/(2-c^{\prime})>1/2$, so the right-hand side would be >1>1, a contradiction. Therefore $c=2$ and $c^{\prime}=0$.
Step 4: conclusion. Let $A:=\{f=2\}$. Then $\mu(A)=1$ and $\nu(A)=0$, hence $\mu\perp\nu$. ∎
Remark 4.
Although the statement of Lemma 3 is reminiscent of classical results based on ergodic decompositions and Choquet theory, the proof given here is entirely independent of such tools. It relies only on invariance, absolute continuity, and the fact that bounded $P$-harmonic functions are constant under ergodicity. In particular, no topological or convex-analytic structure on the set of invariant measures is required.
Proposition 5.
If a Markov kernel $P$ admits at least two distinct invariant probability measures, then it admits two invariant probability measures that are mutually singular.
Proof.
Let $\mu\neq\nu$ be two invariant probability measures.
If both $\mu$ and $\nu$ are ergodic, the result follows directly from Lemma 3.
Otherwise, at least one of them, say $\mu$, is not ergodic. Then there exists a measurable set $A\in\mathcal{E}$ such that
$0<\mu(A)<1\quad\text{and}\quad P(x,A)=\mathbf{1}_{A}(x)\quad\mu\text{-a.s.}$
Define the conditional probability measures
$\mu_{A}(B):=\frac{\mu(B\cap A)}{\mu(A)},\qquad\mu_{A^{c}}(B):=\frac{\mu(B\cap A^{c})}{\mu(A^{c})}.$
Since $A$ is $\mu$-almost surely absorbing, both $\mu_{A}$ and $\mu_{A^{c}}$ are invariant probability measures. Moreover, they are mutually singular since
$\mu_{A}(A)=1\quad\text{and}\quad\mu_{A^{c}}(A)=0.$
This concludes the proof. ∎
Proposition 6.
Let $(E,\mathcal{B})$ be a measurable space and let $P$ be a Markov kernel on $(E,\mathcal{B})$. Assume that $P$ admits a unique invariant probability measure $\mu$. Then $\mu$ is ergodic, i.e. for any $A\in\mathcal{B}$ such that
$P1_{A}=1_{A}\quad\mu\text{-a.s.},$
one has $\mu(A)\in\{0,1\}$.
Proof.
Assume by contradiction that there exists a measurable set $A\in\mathcal{B}$ such that
$P\mathbf{1}_{A}=\mathbf{1}_{A}\quad\mu\text{-a.s.}\qquad\text{and}\qquad 0<\mu(A)<1.$
Define two probability measures $\mu_{1}$ and $\mu_{2}$ on $(E,\mathcal{B})$ by
$\mu_{1}(B)=\frac{\mu(B\cap A)}{\mu(A)},\qquad\mu_{2}(B)=\frac{\mu(B\cap A^{c})}{1-\mu(A)},\qquad B\in\mathcal{B}.$
Since $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-a.s. and $0<\mu(A)<1$, the measures $\mu_{1}$ and $\mu_{2}$ are well defined invariant probability measures and are mutually singular (indeed $\mu_{1}(A)=1$ and $\mu_{2}(A)=0$). In particular, $\mu_{1}\neq\mu_{2}$, which contradicts the uniqueness assumption. Therefore $\mu(A)\in\{0,1\}$ for every $A\in\mathcal{B}$ such that $P\mathbf{1}_{A}=\mathbf{1}_{A}$ $\mu$-almost surely, and $\mu$ is ergodic. ∎
Remark 7.
Proposition 6 shows that ergodicity does not require any additional assumption beyond uniqueness. In particular, no ergodic decomposition or convex-analytic argument is needed to identify the invariant measure as ergodic in the uniqueness regime.
We can now state and prove the main result.
Theorem 8 (Uniqueness via indecomposability).
Let $P$ be a Markov kernel on a standard Borel space $(E,\mathcal{B})$. If $P$ is indecomposable, then $P$ admits at most one invariant probability measure. If one exists, it is ergodic.
Proof.
Assume by contradiction that $P$ admits two distinct invariant probability measures. By Proposition 5, there exist two mutually singular invariant probability measures $\mu_{1}$ and $\mu_{2}$. Hence there exists $A\in\mathcal{B}(E)$ such that
$\mu_{1}(A)=1,\qquad\mu_{2}(A)=0.$
By invariance of $\mu_{1}$ and $\mu_{2}$,
$\mu_{i}(A)=\int_{E}P(x,A)\,\mu_{i}(dx),\qquad i=1,2.$
Since $0\leq P(x,A)\leq 1$, it follows that
$P(x,A)=1\ \ \mu_{1}\text{-a.s.},\qquad P(x,A)=0\ \ \mu_{2}\text{-a.s.}$
Define
$B_{1}:=\bigcap_{n\geq 0}\{x\in E:\;P^{n}(x,A)=1\}.$
For every $n\geq 0$, invariance yields
$\int_{E}P^{n}(x,A)\,\mu_{1}(dx)=\mu_{1}(A)=1,\qquad\int_{E}P^{n}(x,A)\,\mu_{2}(dx)=\mu_{2}(A)=0,$
and therefore
$\mu_{1}\bigl(\{x:\,P^{n}(x,A)=1\}\bigr)=1,\qquad\mu_{2}\bigl(\{x:\,P^{n}(x,A)=1\}\bigr)=0.$
Taking the intersection over $n\geq 0$ gives
$\mu_{1}(B_{1})=1,\qquad\mu_{2}(B_{1})=0.$
Let $x\in B_{1}$. For every $n\geq 0$,
$1=P^{n+1}(x,A)=\int_{E}P^{n}(y,A)\,P(x,dy).$
Since $0\leq P^{n}(y,A)\leq 1$, this implies
$P\bigl(x,\{y:\,P^{n}(y,A)=1\}\bigr)=1\quad\text{for all }n\geq 0.$
Taking the intersection over $n\geq 0$, we obtain
$P(x,B_{1})=1\qquad\text{for all }x\in B_{1}.$
Define
$B_{2}:=\bigcap_{n\geq 0}\{x\in E:\;P^{n}(x,A^{c})=1\}.$
For every $n\geq 0$, invariance yields
$\int_{E}P^{n}(x,A^{c})\,\mu_{2}(dx)=\mu_{2}(A^{c})=1,\qquad\int_{E}P^{n}(x,A^{c})\,\mu_{1}(dx)=\mu_{1}(A^{c})=0,$
and therefore
$\mu_{2}\bigl(\{x:\,P^{n}(x,A^{c})=1\}\bigr)=1,\qquad\mu_{1}\bigl(\{x:\,P^{n}(x,A^{c})=1\}\bigr)=0.$
Taking the intersection over $n\geq 0$ gives
$\mu_{2}(B_{2})=1,\qquad\mu_{1}(B_{2})=0.$
Let $x\in B_{2}$. For every $n\geq 0$,
$1=P^{n+1}(x,A^{c})=\int_{E}P^{n}(y,A^{c})\,P(x,dy).$
Since $0\leq P^{n}(y,A^{c})\leq 1$, this implies
$P\bigl(x,\{y:\,P^{n}(y,A^{c})=1\}\bigr)=1\quad\text{for all }n\geq 0.$
Taking the intersection over $n\geq 0$, we obtain
$P(x,B_{2})=1\qquad\text{for all }x\in B_{2}.$
Thus $B_{1}$ and $B_{2}$ are disjoint nontrivial absorbing measurable sets, contradicting the indecomposability of $P$. Uniqueness follows from Proposition 6. ∎
Remark 9.
On a standard Borel space, $\phi$–irreducibility in the sense of Meyn–Tweedie is a strictly stronger property than indecomposability, since $\phi$–irreducibility rules out the existence of nontrivial absorbing measurable sets. In general, the converse implication does not hold, as indecomposability alone does not preclude purely transient behavior. However, if an invariant probability measure exists, the structural decomposition theory of [9, Section 4.2] implies that indecomposability excludes the presence of more than one closed communicating class and therefore enforces irreducibility with respect to a maximal irreducibility measure $\psi$ in the sense of Meyn–Tweedie. In this sense, indecomposability and $\phi$–irreducibility become equivalent once the existence of an invariant probability measure is guaranteed. From this perspective, indecomposability identifies a purely structural obstruction to uniqueness of invariant probability measures, while stronger assumptions—such as the existence of small or petite sets or additional regularity of the transition kernel—are only required to address recurrence and ergodic convergence properties.
2.1 Quasi–Feller regularity
We now recall the notion of quasi–Feller regularity introduced in [2]. This notion provides a structural framework allowing one to handle transition kernels that are not Feller by factoring the dynamics through a Feller (or strong Feller) kernel on an auxiliary space.
Definition 10 (Quasi–Feller and Quasi–strong Feller, after [2]).
Let $E$ be a Polish space and let $P$ be a Markov transition kernel on $E$. The kernel $P$ is said to be quasi–Feller if there exist
  • a Polish space $W$,
  • a Borel measurable mapping $H:E\to W$ such that $H(K)$ is compact in $W$ for every compact set $K\subset E$,
  • a Markov transition kernel $Q:W\times\mathcal{B}(E)\to[0,1]$
satisfying the following properties:
  1. (i)
    (Feller property of $Q$) For every $f\in C_{b}(E)$, the function
    $Qf:w\mapsto\int_{E}f(y)\,Q(w,dy)$
    belongs to $C_{b}(W)$.
  2. (ii)
    (Factorization) For every bounded measurable function $f$ on $E$ and every $x\in E$,
    $Pf(x)=Qf(H(x)),$
    equivalently,
    $P(x,\cdot)=Q(H(x),\cdot).$
  3. (iii)
    (Essential continuity of $H$) For every invariant probability measure $\mu$ of $P$, one has
    $\mu(D_{H})=0,$
    where $D_{H}$ denotes the set of discontinuity points of $H$.
The kernel $P$ is said to be quasi–strong Feller if, in addition, the kernel $Q$ is strong Feller, i.e. for every bounded measurable function $f$ on $E$, the function $Qf$ is continuous on $W$.
This framework strictly generalizes the classical Feller and strong Feller settings and naturally arises in many non–Feller models with discontinuous dynamics.
We emphasize that the following theorem provides an independent existence and invariance statement. Although it is not required for the subsequent stability results, it isolates a measure–theoretic regularity condition underlying the quasi–Feller framework.
Theorem 11 (Existence and invariance under tightness of iterated kernels).
Let $E$ be a Polish space and let $P$ be a Markov transition kernel on $E$. Fix $x\in E$ and define the averaged iterates
$\nu_{n}^{x}:=\frac{1}{n}\sum_{k=0}^{n-1}P^{k}(x,\cdot),\qquad n\geq 1.$
Assume that:
  1. 1.
    the sequence $(\nu_{n}^{x})_{n\geq 1}$ is tight in $\mathcal{P}(E)$;
  2. 2.
    (essential regularity along limit measures) for every weak limit point $\mu$ of $(\nu_{n}^{x})_{n\geq 1}$ and every $f\in C_{b}(E)$, there exists a bounded Borel function $g:E\to\mathbb{R}$ such that
    $g=Pf\quad\mu\text{--a.s.}\qquad\text{and}\qquad\mu(D_{g})=0,$
    where $D_{g}$ denotes the set of discontinuity points of $g$.
Then $(\nu_{n}^{x})$ admits at least one weak limit point, and every such limit point $\mu$ is an invariant probability measure for $P$. In particular, $P$ admits at least one invariant probability measure.
Remark 12.
This condition should be understood as an a priori version of the essential quasi–Feller principle, stated along limit measures that are not assumed to be invariant.
Proof.
By tightness, there exists a subsequence $(n_{j})_{j\geq 1}$ such that $\nu_{n_{j}}^{x}\Rightarrow\mu$ for some probability measure $\mu$ on $E$. Fix $f\in C_{b}(E)$ and let $g$ be given by assumption (2).
Since $g$ is bounded Borel and $\mu(D_{g})=0$, the portmanteau theorem yields
$\int g\,d\nu_{n_{j}}^{x}\xrightarrow[j\to\infty]{}\int g\,d\mu.$
Since $f$ is continuous, we also have
$\int f\,d\nu_{n_{j}}^{x}\xrightarrow[j\to\infty]{}\int f\,d\mu.$
On the other hand, for all $n\geq 1$,
$\int Pf\,d\nu_{n}^{x}-\int f\,d\nu_{n}^{x}=\frac{1}{n}\bigl(P^{n}f(x)-f(x)\bigr),$
which converges to 0 as $n\to\infty$ since $f$ is bounded. Passing to the subsequence $(n_{j})$ yields
$\lim_{j\to\infty}\int Pf\,d\nu_{n_{j}}^{x}=\lim_{j\to\infty}\int f\,d\nu_{n_{j}}^{x}=\int f\,d\mu.$
Finally, since $g=Pf$ $\mu$–a.s., we have $\int g\,d\mu=\int Pf\,d\mu$, and therefore
$\int Pf\,d\mu=\int f\,d\mu\qquad\text{for all }f\in C_{b}(E).$
This proves that $\mu$ is invariant for $P$. ∎
The following example, adapted from Meyn–Tweedie, shows that the regularity condition introduced in Theorem 11 may hold even when the quasi–Feller property fails.
Example 13 (Essential regularity without quasi-Feller).
Let $E=\mathbb{R}$, let $\varepsilon\in(0,1)$, and let $\nu$ be a probability measure absolutely continuous with respect to Lebesgue measure. Fix two distinct irrational numbers $\alpha\neq\beta$, and define
$T(x)=\begin{cases}\alpha,&x\in\mathbb{Q},\\ \beta,&x\notin\mathbb{Q}.\end{cases}$
Note that $T$ is discontinuous at every point, hence $D_{T}=\mathbb{R}$. Define the Markov kernel
$P(x,A)=(1-\varepsilon)\,\mathbf{1}_{A}(T(x))+\varepsilon\,\nu(A),\qquad x\in E,\ A\in\mathcal{B}(E).$
We first observe that the quasi-Feller property fails. Indeed, since $D_{T}=\mathbb{R}$, we have $\mu(D_{T})=1$ for any probability measure $\mu$. Therefore the condition $\mu(D_{H})=0$ required in Definition 10 cannot be satisfied for any factorization involving $T$, and $P$ is not quasi-Feller.
We now verify the essential regularity condition of Theorem 11. Let $\mu$ be any weak limit point of the empirical measures $\nu_{n}^{x}=\frac{1}{n}\sum_{k=1}^{n}P^{k}(x,\cdot)$. By Theorem 11, $\mu$ is an invariant probability measure for $P$. For every $x\in E$,
$P(x,\mathbb{Q})=(1-\varepsilon)\mathbf{1}_{\mathbb{Q}}(T(x))+\varepsilon\nu(\mathbb{Q})=0,$
since $T(x)\in\{\alpha,\beta\}\subset\mathbb{R}\setminus\mathbb{Q}$ and $\nu(\mathbb{Q})=0$. By invariance of $\mu$,
$\mu(\mathbb{Q})=\int_{E}P(x,\mathbb{Q})\,\mu(dx)=0.$
Hence $T(x)=\beta$ holds $\mu$-almost surely. For any $f\in C_{b}(E)$, it follows that
$Pf(x)=(1-\varepsilon)f(T(x))+\varepsilon\int f\,d\nu=(1-\varepsilon)f(\beta)+\varepsilon\int f\,d\nu,\quad\mu\text{-a.s.}$
Thus $Pf$ is $\mu$-almost surely equal to the constant function
$g\equiv(1-\varepsilon)f(\beta)+\varepsilon\int f\,d\nu,$
which is continuous and satisfies $\mu(D_{g})=0$. This proves that the essential regularity condition of Theorem 11 holds, while the quasi-Feller property fails.
2.2 Almost sure tightness of trajectories
As a preparatory step, we recall a classical criterion ensuring almost sure tightness of empirical occupation measures. In the quasi–Feller framework, long–time control of the trajectories is typically obtained through Lyapunov–Hájek type drift conditions, which imply that the Markov chain spends an asymptotically negligible proportion of time outside compact subsets of the state space.
More precisely, under suitable Lyapunov assumptions, the empirical occupation measures
$\Lambda_{n}(\omega,\cdot)\;:=\;\frac{1}{n}\sum_{k=0}^{n-1}\delta_{X_{k}(\omega)}(\cdot)$
are almost surely tight under $P_{x}$, for every initial condition $x\in E$.
This result is classical and follows from a standard Lyapunov–martingale argument; see, for instance, [7, Appendix H] or [1]. Since this argument is orthogonal to the main contribution of the present paper, we do not reproduce the proof here.
2.3 Stability via uniqueness
We now combine the tightness property recalled above with the uniqueness result obtained in Section 2. Recall that the notion of stability was introduced by Duflo [6] and refers to almost sure convergence of time averages for bounded continuous observables.
Definition 14 (Stability in the sense of Duflo).
A Markov chain $(X_{n})_{n\geq 0}$ is said to be stable if there exists a probability measure $\mu$ such that, for every $x\in E$ and every $f\in C_{b}(E)$,
$\frac{1}{n}\sum_{k=0}^{n-1}f(X_{k})\;\xrightarrow[n\to\infty]{P_{x}\text{-a.s.}}\;\int f\,d\mu.$
Remark 15.
Stability does not imply that the limiting measure $\mu$ is invariant. Counterexamples can be found in [1]. As shown below, quasi–Feller regularity provides a sufficient structural condition ensuring invariance, consistently with the essential quasi–Feller principle stated above.
Theorem 16 (Stability induced by uniqueness).
Let $P$ be a quasi–Feller Markov kernel such that, for every $x\in E$, the empirical occupation measures $(\Lambda_{n})$ are $P_{x}$–almost surely tight. Under these assumptions, $P$ admits at least one invariant probability measure. If this invariant probability measure is unique, then the Markov chain $(X_{n})_{n\geq 0}$ is stable in the sense of Duflo.
Proof.
The result follows from Theorem H.2 in [1]. ∎
2.4 From Duflo stability to positive Harris recurrence
We now explain how a strengthened version of the essential regularity condition allows one to extend Duflo stability from bounded continuous functions to bounded Borel functions, without invoking any structural quasi–strong Feller factorization.
Proposition 17 (Ergodic averages under essential strong regularity).
Let $(X_{n})_{n\geq 0}$ be a Markov chain on a Polish space $E$ with transition kernel $P$. Assume that:
  1. 1.
    the chain is stable in the sense of Duflo;
  2. 2.
    $\mu$ is an invariant probability measure for $P$;
  3. 3.
    (essential strong regularity) for every bounded Borel function $f\in\mathcal{B}_{b}(E)$, there exists a bounded continuous function $g\in C_{b}(E)$ such that
    $g=Pf\quad\mu\text{--a.s.}\qquad\text{and}\qquad\mu(D_{g})=0.$
Then, for every bounded Borel function $f\in\mathcal{B}_{b}(E)$ and every $x\in E$,
$\frac{1}{n}\sum_{k=0}^{n-1}f(X_{k})\xrightarrow[n\to\infty]{P_{x}\text{-a.s.}}\int f\,d\mu.$
2.5 Harris-recurrence of Essential Feller Transition Kernel
Under Quasi-Feller regularity and assuming that the support of a maximal irreducibility measure $\psi$ has nonempty interior—together with classical assumptions ensuring tightness from all initial conditions—positive Harris recurrence is established in [2] without assuming the existence of a petite set. Rather, under these hypotheses, the existence of a petite set is derived as an intermediate result, which then allows one to re-enter the classical Meyn–Tweedie framework a posteriori.
In the present setting, essential Feller regularity and tightness ensure the existence of an invariant probability measure $\mu$. Moreover, the indecomposability assumption rules out the presence of disjoint absorbing components and therefore implies $\psi$-irreducibility for a maximal irreducibility measure $\psi$ in the sense of Meyn–Tweedie. In this context, the assumption that $\operatorname{supp}(\psi)$ has nonempty interior plays exactly the same role as in [2]: combined with essential Feller regularity, it allows one to recover the existence of a petite set, and hence to establish Harris recurrence within the Meyn–Tweedie framework.
Remark. At this stage, the role of the Meyn–Tweedie framework is to provide an accessibility and regeneration mechanism through which stability properties can be upgraded into recurrence properties. It is therefore natural to ask whether positive Harris recurrence alone already enforces the essential regularity condition introduced above.
Remark. Beyond its conceptual interest, the essential regularity framework appears to cover a broad class of examples encountered in practice. Indeed, many Markov kernels that fail to satisfy classical Feller or quasi–Feller assumptions arise from the combination of a highly irregular deterministic component with a smoothing noise term. In such situations, although standard regularity properties may be violated, the transition kernel often satisfies the essential regularity condition along invariant or limit measures. The following example shows that this conclusion is not universal, even under a strong Meyn–Tweedie-type minorization condition.
Example 18 (Positive Harris recurrence without essential regularity).
Let $E=[0,1]$ endowed with its Borel $\sigma$–field and let $\nu$ denote the Lebesgue probability measure on $E$. Fix $\varepsilon\in(0,1)$. Let $C\subset E$ be a closed set with empty interior and positive Lebesgue measure (for instance a fat Cantor set). Define the measurable mapping $T:E\to\{0,1\}$ by
$T(x)=\begin{cases}0,&x\in C,\\ 1,&x\in C^{c},\end{cases}$
and define a Markov transition kernel $P$ on $E$ by
$P(x,A)=(1-\varepsilon)\,\mathbf{1}_{A}(T(x))+\varepsilon\,\nu(A),\qquad x\in E,\ A\in\mathcal{B}(E).$
Harris recurrence. Since $P(x,\cdot)\geq\varepsilon\nu(\cdot)$ for all $x\in E$, the chain satisfies a uniform Doeblin minorization. In particular, it is $\nu$–irreducible and the whole space $E$ is a small set in the sense of Meyn–Tweedie. As a consequence, the chain is positive Harris recurrent and admits a unique invariant probability measure $\pi$.
Failure of essential regularity. Let $f\in C_{b}(E)$ be such that $f(0)\neq f(1)$. Then
$Pf(x)=(1-\varepsilon)f(T(x))+\varepsilon\int_{E}f\,d\nu=\begin{cases}a,&x\in C,\\ b,&x\in C^{c},\end{cases}\qquad a\neq b.$
Since $C$ is closed with empty interior, one has $\partial C=C$, and therefore every point of $C$ is a point of discontinuity of $Pf$. Moreover, the invariant probability measure $\pi$ satisfies $\pi\geq\varepsilon\nu$, so that $\pi(C)>0$. Consequently,
$\pi(D_{Pf})>0.$
It follows that there exists no bounded Borel function $g$ such that $g=Pf$ $\pi$–almost surely and $\pi(D_{g})=0$. Hence the essential regularity condition fails for this kernel.
Remark 19.
This example shows that positive Harris recurrence does not imply essential regularity, even in a compact state space and under a strong Doeblin-type minorization. It illustrates that the Meyn–Tweedie framework and the essential regularity approach address genuinely distinct aspects of the long–time behavior of Markov chains: the former focuses on recurrence and regeneration properties, while the latter is tailored to ergodic properties of time averages.
3 Further remarks
Remark 20 (A topological interpretation).
In many continuous–state models, the positivity mechanism underlying the uniqueness results of this paper can be verified through topological support properties of the transition kernel, providing concrete sufficient conditions for indecomposability.
For instance, assume that there exists a closed set $F\subset E$ such that every invariant probability measure is supported on $F$, and that for all $x\in F$ there exists $n\geq 1$ such that the support of $P^{n}(x,\cdot)$ has nonempty interior relative to $F$. Then any $\sigma$–finite measure $\psi$ charging open subsets of $F$ is necessarily charged by all invariant probability measures. Since distinct invariant ergodic measures are mutually singular, uniqueness follows from the impossibility for two such measures to both charge the same reference measure.
Such arguments show that indecomposability, and hence uniqueness, may be established from topological considerations alone, without requiring full $\phi$–irreducibility on the whole state space.
Remark 21 (Uniqueness without recurrence).
The arguments developed in this paper do not rely on any recurrence assumption for the original Markov kernel. In particular, uniqueness may hold whenever an invariant probability measure exists, even if the chain is not positive recurrent.
This contrasts with classical approaches based on Harris recurrence, where uniqueness is typically obtained together with strong ergodic properties. The present results show that these notions can be separated: recurrence is a dynamical property governing long–time returns, while uniqueness emerges here as a purely structural consequence of indecomposability, revealed through the one–step positivity mechanism induced by the resolvent kernel.
Remark 22 (On geometric convergence under additional contraction assumptions).
The present work deliberately focuses on existence and uniqueness of invariant probability measures, without addressing quantitative rates of convergence. Nevertheless, stronger ergodic properties can be recovered under additional contraction assumptions.
For instance, consider a Markov chain of the form
$X_{n+1}=f(X_{n})+\varepsilon_{n+1},$
where $(\varepsilon_{n})_{n\geq 1}$ are i.i.d. random variables whose law admits a density with respect to Lebesgue measure. Assume that the chain satisfies a Lyapunov drift condition
$PV\leq\alpha V+b,\qquad\alpha<1,$
with relatively compact sublevel sets, and that there exist $R>0$ and $\rho<1$ such that
$V(x),V(y)\leq R\quad\Longrightarrow\quad W_{d}\bigl(P(x,\cdot),P(y,\cdot)\bigr)\leq\rho\,d(x,y),$
for some bounded distance $d$.
Then the chain converges geometrically fast to its invariant distribution in a Wasserstein distance associated with $d$ (possibly after a suitable Lyapunov weighting). This illustrates that geometric ergodicity relies on a genuine contraction property, which is logically independent of the uniqueness mechanism isolated in the present work.
Remark 23 (Relation with ergodic decomposition).
Invariant probability measures admit a decomposition into ergodic components, and distinct invariant ergodic measures are mutually singular. The contribution of the present work is to show that, under the positivity mechanism induced by the resolvent kernel, all invariant probability measures are forced to charge a common $\sigma$–finite reference measure.
This structural constraint prevents the coexistence of several ergodic components and therefore yields uniqueness. From this perspective, the resolvent kernel acts as a device that transforms finite–time reachability into a one–step positivity constraint at the level of invariant measures.
4 Conclusion
The main contribution of this paper is to identify indecomposability as a purely structural condition ensuring uniqueness of invariant probability measures for discrete–time Markov kernels. We show that, on a standard Borel space, indecomposability alone is sufficient to rule out the coexistence of several invariant probability measures, independently of any recurrence, regeneration or ergodic convergence property, provided existence is known.
A key observation is that classical irreducibility assumptions are not essential for uniqueness in themselves, but rather provide convenient sufficient conditions for indecomposability. This allows uniqueness to be separated conceptually from dynamical notions such as Harris recurrence or minorization, which are traditionally used to derive uniqueness together with convergence properties.
We further show that, under irreducibility, indecomposability can be enforced through a resolvent–type kernel. The resolvent transforms finite–time reachability into a one–step positivity property at the level of invariant measures, forcing all invariant probability measures to charge a common $\sigma$–finite reference measure. Since distinct invariant ergodic measures are mutually singular, this excludes the possibility of multiple ergodic components and yields uniqueness without any appeal to return–time estimates or petite set constructions.
From a broader perspective, the present results clarify the respective roles of structural and dynamical assumptions in the theory of invariant measures. Recurrence, regeneration and minorization govern long–term behaviour and ergodic convergence, while uniqueness emerges here as a structural consequence of indecomposability. This viewpoint complements the quasi–Feller approach of [2], which addresses existence and stability for non–Feller models, and leads to a more modular understanding of invariant probability measures for a wide class of discrete–time Markov chains, including discontinuous and non–regular dynamics.